OpenStudy (anonymous):

Solve: -3x - 1 = 8 Answer 6 -3 -2.3 3

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

-3x -1 = 8 So, add -1 to both side; we will get -3x = 9. Then we divide by -3 that will give us x = -3. So, the answer is -3. We can also check our answer by plugging in -3; -3(-3) -1 = 8; 9 - 1 =8 8 =8.

OpenStudy (anonymous):

-3x-1=8 -3x=8+1 -3x=9 x=-9/3 x=-3